# Decoupled Multimodal Fusion for User Interest Modeling in Click-Through Rate Prediction

### v1

@@ Abstract.
Modern industrial recommendation systems improve recommendation performance by integrating multimodal representations from pre-trained models into ID-based Click-Through Rate (CTR) prediction frameworks. However, existing approaches typically adopt modality-centric modeling strategies that process ID-based and multimodal embeddings independently, failing to capture fine-grained interactions between content semantics and behavioral signals. In this paper, we propose Decoupled Multimodal Fusion (DMF), which introduces a modality-enriched modeling strategy to enable fine-grained interactions between ID-based collaborative representations and multimodal representations for user interest modeling. Specifically, we construct target-aware features to bridge the semantic gap across different embedding spaces and leverage them as side information to enhance the effectiveness of user interest modeling. Furthermore, we design an inference-optimized attention mechanism that decouples the computation of target-aware features and ID-based embeddings before the attention layer, thereby alleviating the computational bottleneck introduced by incorporating target-aware features. To achieve comprehensive multimodal integration, DMF combines user interest representations learned under the modality-centric and modality-enriched modeling strategies. Offline experiments on public and industrial datasets demonstrate the effectiveness of DMF. Moreover, DMF has been deployed on the product recommendation system of the international e-commerce platform Lazada, achieving relative improvements of 5.30% in CTCVR and 7.43% in GMV with negligible computational overhead.
Click-through Rate Prediction, User Interest Modeling, Multimodal Fusion, Target-aware Attention
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
1. Introduction
@@
Click-Through Rate (CTR) prediction is a fundamental task in industrial recommender systems, aiming to estimate the probability that a user will click on a candidate item. The predicted CTR serves as a crucial signal for ranking decisions (Guo et al., 2017). User interest modeling, which focuses on capturing evolving user preferences from historical interaction sequences, has become a key component of modern CTR prediction models and has contributed to notable performance improvements (Zhou et al., 2018; Xia et al., 2023). However, most industrial systems still rely solely on sparse ID features, which often lead to insufficient item representation learning, thereby degrading CTR prediction performance. In addition, ID-based embeddings effectively capture behavioral collaboration from user interactions, yet fail to represent the rich semantic information embedded in multimodal content.
@@
Recently, the remarkable success of Large Language Models (LLMs) across various fields has spurred growing interest in integrating multimodal information into recommendation systems, with the goal of leveraging the rich world knowledge encoded in these models (Yuan et al., 2023; Hou et al., 2022). As online computing resources are limited and response time constraints are strict, industrial systems typically adopt a two-stage framework to incorporate multimodal information into ID-based CTR prediction models (Sheng et al., 2024). The framework first extracts multimodal representations from pre-trained models and then integrates them into the downstream ID-based model while keeping the representations frozen. In the integration stage, directly incorporating multimodal representations into ID-based models suffers from the semantic gap across different embedding spaces, leading to suboptimal performance (Pan et al., 2024).
@@
To mitigate the misalignment, recent approaches leverage multimodal similarity relationships between target items and historical interaction sequences as proxies for aligning embedding spaces, resulting in improved recommendation performance (Sheng et al., 2024; Zhang et al., 2025). These approaches can be broadly categorized into two types: histogram-based and attention-based approaches. The histogram-based approach (Sheng et al., 2024) converts high-dimensional multimodal representations into a compact similarity histogram by partitioning pairwise similarity scores into several discrete tiers and counting the number of matches in each bin. The attention-based approach (Zhang et al., 2025) uses encoded similarity embeddings as item representations for behavioral-level multimodal interest fusion. However, both approaches adopt modality-centric modeling strategies that process ID-based and multimodal embeddings independently, failing to capture fine-grained interactions between content semantics and behavioral signals.
@@
When seeking to enable fine-grained interactions between ID-based collaborative representations and multimodal representations, one might naturally consider treating multimodal similarity scores as side information and feeding them together with ID-based embeddings into a user interest modeling module, such as the Target Attention (TA) network (Zhou et al., 2024). Although this modality-enriched modeling approach is straightforward, it faces two challenges when deployed in large-scale industrial recommendation systems: First, multimodal similarity scores are target-aware features, which means that different candidate items within the same request yield distinct similarity score sequences, even when derived from an identical historical interaction sequence. Using such features as side information requires recomputing the key ($K$) and value ($V$) representations for each candidate during online inference, making it impossible to precompute or cache $K$ and $V$ (Chang et al., 2023). The prohibitive computational overhead is unacceptable for industrial recommendation systems with more than $10^{3}$ candidate items (Chen et al., 2021; Cao et al., 2022). Second, even if computationally feasible, relying solely on fusing ID-based collaborative representations and multimodal representations in user interest modeling may lead to underutilization of multimodal information, as behavioral-level fusion is insufficient to capture a holistic representation of user multimodal interests.
@@
To address these challenges, we propose Decoupled Multimodal Fusion (DMF), a hybrid framework that introduces a modality-enriched modeling strategy to enable fine-grained interactions between ID-based collaborative representations and multimodal representations for user interest modeling, and combines modality-centric and modality-enriched paradigms to leverage their complementary strengths for boosting CTR prediction performance. For the first challenge, we propose Decoupled Target Attention (DTA) to integrate target-aware features into user interest modeling without sacrificing inference efficiency. For the second challenge, to fully exploit multimodal signals, we propose the Complementary Modality Modeling (CMM) module, which fuses user interest representations learned under modality-centric and modality-enriched frameworks.
@@
  • We introduce a novel modality-enriched modeling paradigm that enables fine-grained interaction between ID-based and multimodal representations by treating multimodal similarity as target-aware side information, effectively bridging the semantic gap.
@@
  • We design Complementary Modality Modeling (CMM) to unify the strengths of modality-centric and modality-enriched strategies, forming a comprehensive user interest representation. Extensive experiments show that DMF outperforms state-of-the-art methods, and has been successfully deployed on Lazada’s product recommendation system, achieving significant gains in CTCVR (+5.30%) and GMV (+7.43%).
2. Related Work
2.1. Click-through Rate Prediction
@@
CTR prediction plays an important role in modern recommendation systems, aiming to estimate the likelihood of a user clicking a candidate item (Lin et al., 2024; Liu et al., 2024). Approaches like DeepFM (Guo et al., 2017) and DCN (Wang et al., 2017) are designed for modeling interactions. Sequential approaches focus on modeling user interests from historical interaction sequences, which can be roughly classified into target-agnostic approaches and target-aware approaches. Approaches like SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) are formulated in a target-agnostic fashion, where user interaction sequences are independent of the candidate item and the derived user representations remain fixed across different candidate items. In contrast, most mainstream CTR prediction approaches (Chang et al., 2023; Zhou et al., 2018; Cao et al., 2022; Xia et al., 2023; Ma et al., 2024) used in industrial settings are formulated in a target-aware fashion that adaptively models user interests by considering the relevance between candidate items and historical interactions. However, these approaches rely solely on sparse ID features, which fail to represent the rich semantic information embedded in multimodal content, thereby degrading CTR prediction performance.
2.2. Multimodal Recommendation
@@
The rich semantic information embedded in multimodal data has spurred extensive efforts to incorporate it into recommendation systems. The potential of learning multimodal representations in an end-to-end manner alongside recommendation model training has been investigated in prior work (Yuan et al., 2023). Despite their effectiveness, these approaches incur high computational costs, rendering them impractical for large-scale industrial deployment. Recently, a two-stage framework (Pan et al., 2024; Pal et al., 2020) has been introduced in industrial systems to leverage multimodal data for both effectiveness and cost efficiency. Building on this success, some studies have adopted the well-established two-stage paradigm and focus on improving the integration stage (Sheng et al., 2024; Zhang et al., 2025).
@@
Since directly incorporating multimodal representations into ID-based models is ineffective due to distinct spaces (Yang et al., 2025; Cui et al., 2025), a variety of approaches have been proposed to address this misalignment. MARN (Li et al., 2020) draws inspiration from Generative Adversarial Networks (GANs), leveraging an adversarial mechanism to achieve alignment across modalities. MAKE (Sheng et al., 2024) proposes a three-stage framework that decouples the optimization of multimodal related parameters from that of ID features via a recommendation pre-training module. By encoding similarity scores into learnable embeddings, DMAE (Zhang et al., 2025) aligns heterogeneous semantic spaces without requiring direct compatibility between raw embedding vectors. As similarity encoding avoids the instability and complexity of adversarial or multi-stage alignment by distilling relational knowledge from external embeddings into a learnable space, in this paper, we employ multimodal similarity scores between target items and historical interaction sequences as alignment proxies and side information to enhance the effectiveness of user interest modeling.
@@
Within the similarity-based alignment framework, existing multimodal fusion approaches can be broadly categorized into histogram-based modeling and attention-based modeling. Histogram-based modeling (Sheng et al., 2024) calculates a statistical similarity histogram for user interest modeling. Attention-based modeling (Zhang et al., 2025) utilizes the encoded similarity embeddings as input representations for behavioral-level multimodal interest fusion. However, both approaches process ID-based and multimodal embeddings independently, failing to capture fine-grained interactions between content semantics and behavioral signals. This limitation motivates the development of a more expressive multimodal fusion approach.
2.3. Side Information Fusion
@@
In sequential recommendation, side information fusion approaches integrate various item attributes into the overall item representation, enabling the model to capture diverse collaborative signals across items (Kim et al., 2025). From the perspective of when feature representations are merged, existing self-attention-based side information fusion approaches can be classified into early, late, and hybrid fusion (Wang et al., 2024).
@@
Early fusion (Zhou et al., 2020) integrates item ID and side information before feeding them into the model, facilitating interaction between item ID and side information representations. However, this simple aggregation may lead to information invasion (Liu et al., 2021) due to inherent differences in representation spaces. In contrast, late fusion postpones merging item ID and side information until the final prediction stage. For example, FDSA (Zhang et al., 2019) uses independent self-attention blocks to model item ID and side information in isolation, thereby preserving their distinct contextual patterns. Although such isolation supports accurate modeling within each modality, it inherently limits the ability to learn fine-grained interactions between item ID and side information. Hybrid fusion lies between early and late fusion, allowing item ID and side information to interact in the middle layer. NOVA (Liu et al., 2021) proposes only to incorporate attributes in the calculation of attention scores to mitigate the information invasion problem. DIF-SR (Xie et al., 2022) decouples the attention calculation of various side information and item ID representation, allowing higher-rank attention matrices and flexible gradients. ASIF (Wang et al., 2024) utilizes side information without noisy interference via fused attention with untied position information.
@@
3. Preliminaries
Figure 1. The framework of DMF. Multimodal embeddings are used to compute similarity scores between the target item and each historical interaction. These scores are mapped into learnable embeddings via the MSE module for input to the DTA module, while the Similarity Histogram operates directly on the original scores. The resulting interest representations are fused by the CMM module and combined with outputs from auxiliary tasks for CTR prediction.:
3.1. Task Formulation
@@
The goal of CTR prediction task is to estimate the probability that a user will click on a candidate item. Formally, let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{|\mathcal{D}|}$ denote the training set, where $x_{i}$ represents the input features of the $i$-th instance, and $y_{i}\in\{0,1\}$ is the binary label indicating whether a click occurred ($y_{i}=1$) or not ($y_{i}=0$). The predicted CTR is computed as:
(1) $\hat{y}_{i}=\mathcal{F}\left(\mathcal{E}\left(x_{i}\right)\right),$
@@
where $\mathcal{F}(\cdot)$ is the prediction model and $\mathcal{E}(\cdot)$ is an embedding layer that is adopted to transform the input $x_{i}$ into dense embeddings.
Given $x_{i}$, the CTR model is trained to predict $\hat{y}_{i}$ by minimizing the cross-entropy loss:
(2) $\mathcal{L}=-\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\left(y_{i}\log\hat{y}_{i}+\left(1-y_{i}\right)\log\left(1-\hat{y}_{i}\right)\right).$
3.2. Target-aware Attention
Effective user interest modeling models are usually based on target-aware attention, which adaptively captures user interests from historical interactions concerning the target item (Zhou et al., 2018; Chang et al., 2023; Cao et al., 2022; Zhou et al., 2024). In CTR prediction tasks, each item in a given interaction sequence serves as both a key (K) and a value (V), while the target item acts as the query (Q).
Formally, let $Q\in\mathbb{R}^{B\times d}$ denote the embedding matrix of $B$ candidate items, and $S\in\mathbb{R}^{B\times L\times d}$ represent the user behavior sequence, where $L$ is the maximum length of historical interactions and $d$ is the hidden dimension. For a specific target item $q_{i}\in\mathbb{R}^{d}$, the attention weights over the sequence are computed via scaled dot-product similarity:
(3) $\alpha_{ij}=\frac{\exp(q_{i}^{\top}s_{j}/\tau)}{\sum_{k=1}^{L}\exp(q_{i}^{\top}s_{k}/\tau)},$
where $\tau$ denotes a temperature scaling factor that stabilizes the softmax distribution by preventing excessively large logits (Vaswani et al., 2017). The resulting user interest representation is then obtained as a weighted sum of the sequence elements:
(4) $\text{Attention}(q_{i},S)=\sum_{j=1}^{L}\alpha_{ij}s_{j}.$
DIN (Zhou et al., 2018) and MHTA (Chang et al., 2023) are the two most representative target-aware attention approaches. The main difference between them is how the attention weight $\alpha_{ij}$ is obtained. In DIN, the relevance score is derived through a deep interaction network: the embedding of each behavior item is concatenated or element-wise interacted with that of the target item, followed by a multilayer perceptron (MLP) to produce an attention weight. This fully learnable, non-linear interaction captures complex feature crosses but requires recomputing the entire attention module for every candidate item during inference. In contrast, MHTA computes attention weights via scaled dot-product operations after conducting linear projection on Q, K, and V. In practical, the linear projections of K and V in MHTA attention computation form the primary computational bottleneck, whereas the main computational cost in DIN arises from the MLP.
4. Methodology

### v2

@@ Abstract
Modern industrial recommendation systems enhance Click-Through Rate (CTR) prediction by integrating frozen multimodal representations from pre-trained models into ID-based frameworks. However, existing approaches predominantly adopt modality-centric modeling strategies that process ID and multimodal signals in isolation, thereby failing to capture fine-grained interactions between behavioral dynamics and content semantics. In this paper, we propose Decoupled Multimodal Fusion (DMF), a novel framework that introduces modality-enriched modeling to enable dynamic, context-aware fusion of ID-based and multimodal representations during user interest encoding. We design Decoupled Target Attention (DTA), which avoids redundant computation by reusing target-agnostic ID-derived components and integrating similarity signals via discretized embedding lookup. Furthermore, DMF unifies modality-centric and modality-enriched pathways through Complementary Modality Modeling (CMM) to balance semantic generalization and behavioral personalization. Offline experiments on public and industrial datasets demonstrate the effectiveness of DMF. More importantly, under industrial-scale serving conditions with 1200 candidate items, DTA achieves 3× higher QPS than early fusion, enabling high-throughput deployment with negligible latency overhead. DMF has been successfully deployed on Lazada’s product recommendation system, yielding significant improvements of +5.30% in CTCVR and +7.43% in GMV.
\$*\$\$*\$footnotetext: Corresponding author.
I Introduction
@@
Click-Through Rate (CTR) prediction is a fundamental task in industrial recommender systems, aiming to estimate the probability that a user will click on a candidate item. The predicted CTR serves as a crucial signal for ranking decisions [1, 2, 3, 4]. User interest modeling, which focuses on capturing evolving user preferences from historical interaction sequences, has become a key component of modern CTR prediction models and has contributed to notable performance improvements [5, 6]. However, most industrial systems still rely solely on sparse ID-based features, which often lead to insufficient item representation learning, thereby degrading CTR prediction performance. In addition, ID-based embeddings effectively capture behavioral collaboration from user interactions, yet fail to represent the rich semantic information embedded in multimodal content.
@@
Recent advances in multimodal foundation models have inspired efforts to inject semantic knowledge into recommendation systems [7, 8]. As online computing resources are limited and response time constraints are strict, industrial systems typically adopt a two-stage framework to incorporate multimodal information into ID-based CTR prediction models [9]. The framework first extracts multimodal representations from pre-trained models and then integrates them into the downstream ID-based model while keeping the representations frozen. In the integration stage, directly incorporating multimodal representations into ID-based models suffers from the semantic gap across different embedding spaces, leading to suboptimal performance [10].
@@
To mitigate the misalignment, recent approaches leverage multimodal similarity relationships between target items and historical interaction sequences as proxies for aligning embedding spaces, resulting in improved recommendation performance [9, 11]. The similarity histogram-based approach [9] converts high-dimensional multimodal representations into a compact similarity histogram by partitioning pairwise similarity scores into several discrete tiers and counting the number of matches in each bucket. The attention-based approach [11] uses encoded similarity embeddings as item representations for behavioral-level multimodal interest fusion. However, both approaches adopt modality-centric modeling strategies (Fig. 1(a)) that process ID-based embeddings and multimodal representations independently, falling short in capturing fine-grained interactions between content semantics and behavioral signals.
Figure 1: (a) Modality-centric Modeling: ID-based embeddings and multimodal representations are encoded independently, without fine-grained interaction during user interest modeling. (b) Modality-enriched Modeling: Multimodal representations are integrated as side information into the interest encoder, enabling fine-grained interaction between content semantics and behavioral signals during user interest modeling. (c) Complementary Modality Modeling (CMM): A structured fusion strategy that combines modality-centric and modality-enriched pathways. This hybrid approach produces a comprehensive user interest representation balancing semantic generalization and fine-grained behavioral personalization.
@@
When seeking to enable fine-grained interactions between ID-based collaborative representations and multimodal representations, one might naturally consider treating multimodal similarity scores as side information and feeding them together with ID-based embeddings into a user interest modeling module, such as the Target Attention (TA) network [12]. Although this modality-enriched modeling (Fig. 1(b)) approach is straightforward, it faces two challenges when deployed in large-scale industrial recommendation systems: First, multimodal similarity scores are target-aware features, which means that different candidate items within the same request yield distinct similarity score sequences, even when derived from an identical historical interaction sequence. Using such features as side information requires recomputing the key ($K$) and value ($V$) representations for each candidate during online inference, making it impossible to reuse $K$ and $V$ [13]. The prohibitive computational overhead is unacceptable for industrial recommendation systems with more than $10^{3}$ candidate items [14, 15]. Second, even if computationally feasible, relying solely on fusing ID-based collaborative representations and multimodal representations in user interest modeling may lead to underutilization of multimodal information, as behavioral-level fusion is insufficient to capture a holistic representation of user multimodal interests.
@@
To address these challenges, we propose Decoupled Multimodal Fusion (DMF), a hybrid framework that introduces a modality-enriched modeling strategy to enable fine-grained interactions between ID-based collaborative representations and multimodal representations for user interest modeling, and combines modality-centric and modality-enriched paradigms to leverage their complementary strengths for boosting CTR prediction performance. For the first challenge, we propose Decoupled Target Attention (DTA) to integrate target-aware features into user interest modeling without sacrificing inference efficiency. For the second challenge, to fully exploit multimodal signals, we propose Complementary Modality Modeling (CMM)(Fig. 1(c)), which fuses user interest representations learned under modality-centric and modality-enriched frameworks.
@@
  • We identify the limitations of existing modality-centric approaches in capturing fine-grained semantics and propose modality-enriched modeling as a new paradigm that dynamically integrates multimodal signals into behavioral sequence encoding.
@@
  • We design Complementary Modality Modeling (CMM) to unify the strengths of modality-centric modeling and modality-enriched modeling strategies, achieving both semantic generalization and behavioral personalization. Extensive experiments demonstrate that DMF outperforms state-of-the-art methods and has been successfully deployed in the product recommendation system of the international e-commerce platform Lazada, yielding significant gains in CTCVR (+5.30%) and GMV (+7.43%).
II Related Work
In this section, we review related work in three major areas, i.e., click-through rate prediction, multimodal recommendation, and side information fusion.
II-A Click-through Rate Prediction
@@
CTR prediction plays an important role in modern recommendation systems, aiming to estimate the likelihood of a user clicking a candidate item [16, 17]. Approaches like DeepFM [1] and DCN [18] are designed for modeling interactions. Sequential approaches focus on modeling user interests from historical interaction sequences, which can be roughly classified into target-agnostic approaches and target-aware approaches. Approaches like SASRec [19] and BERT4Rec [20] are formulated in a target-agnostic fashion, where user interaction sequences are independent of the candidate item and the derived user representations remain fixed across different candidate items. In contrast, most mainstream CTR prediction approaches [13, 5, 15, 6, 21] used in industrial settings are formulated in a target-aware fashion that adaptively models user interests by considering the relevance between candidate items and historically interacted items. However, these approaches rely solely on sparse ID-based features, which fail to represent the rich semantic information embedded in multimodal content, thereby degrading CTR prediction performance.
II-B Multimodal Recommendation
@@
The rich semantic information embedded in multimodal data has spurred extensive efforts to incorporate it into recommendation systems. The potential of learning multimodal representations in an end-to-end manner alongside recommendation model training has been investigated in prior work [7]. Despite their effectiveness, these approaches incur high computational costs, rendering them impractical for large-scale industrial deployment. Recently, a two-stage framework [10, 22] has been introduced in industrial systems to leverage multimodal data for both effectiveness and cost efficiency. Building on this success, some studies have adopted the well-established two-stage paradigm and focus on improving the integration stage [9, 11].
@@
Since directly incorporating multimodal representations into ID-based models is ineffective due to distinct spaces [23, 24], a variety of approaches have been proposed to address this misalignment. MARN [25] draws inspiration from Generative Adversarial Networks (GANs), leveraging an adversarial mechanism to achieve alignment across modalities. MAKE [9] proposes a three-stage framework that decouples the optimization of multimodal related parameters from that of ID-based features via a recommendation pre-training module. By encoding similarity scores into learnable embeddings, DMAE [11] aligns heterogeneous semantic spaces without requiring direct compatibility between raw embedding vectors. As similarity encoding avoids the instability and complexity of adversarial or multi-stage alignment by distilling relational knowledge from external embeddings into a learnable space, in this paper, we employ multimodal similarity scores between target items and historically interacted items as alignment proxies and side information to enhance the effectiveness of user interest modeling.
@@
Within the similarity-based alignment framework, existing multimodal fusion approaches can be broadly categorized into similarity histogram-based modeling and attention-based modeling. Similarity histogram-based modeling [9] calculates a statistical similarity histogram for user interest modeling. Attention-based modeling [11] utilizes the encoded similarity embeddings as input representations for behavioral-level multimodal interest fusion. However, both approaches process ID-based embeddings and multimodal representations independently, failing to capture fine-grained interactions between content semantics and behavioral signals. This limitation motivates the development of a more expressive multimodal fusion approach.
II-C Side Information Fusion
@@
In sequential recommendation, side information fusion approaches integrate various item attributes into the overall item representation, enabling the model to capture diverse collaborative signals across items [26]. From the perspective of when feature representations are merged, existing self-attention-based side information fusion approaches can be classified into early, late, and hybrid fusion [27].
@@
Early fusion [28] integrates item ID and side information before feeding them into the model, facilitating interaction between item ID and side information representations. However, this simple aggregation may lead to information invasion [29] due to inherent differences in representation spaces. In contrast, late fusion postpones merging item ID and side information until the final prediction stage. For example, FDSA [30] uses independent self-attention blocks to model item ID and side information in isolation, thereby preserving their distinct contextual patterns. Although such isolation supports accurate modeling within each modality, it inherently limits the ability to learn fine-grained interactions between item ID and side information. Hybrid fusion lies between early and late fusion, allowing item ID and side information to interact in the middle layer. NOVA [29] proposes only to incorporate attributes in the calculation of attention scores to mitigate the information invasion problem. DIF-SR [31] decouples the attention calculation of various side information and item ID representation, allowing higher-rank attention matrices and flexible gradients. ASIF [27] utilizes side information without noisy interference via fused attention with untied position information.
@@
III Preliminaries
III-A Task Formulation
@@
The goal of CTR prediction is to estimate the probability that a user clicks a candidate item, referred to as the target item. Formally, let $\mathcal{D}=\{(x_{i},y_{i})\}$ denote the training set, where $x_{i}$ represents the input feature vector of the $i$-th instance and $y_{i}\in\{0,1\}$ is the binary label indicating whether a click occurs ($y_{i}=1$) or not ($y_{i}=0$). The predicted CTR is computed as:
$\hat{y}_{i}=\mathcal{F}(x_{i}),$ (1)
@@
where $\mathcal{F}(\cdot)$ denotes the CTR prediction model, which includes an embedding layer mapping categorical features into dense vectors.
The model is trained by minimizing the binary cross entropy loss:
$\mathcal{L}=-\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\left(y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})\right).$ (2)
Figure 2: The framework of DMF. Multimodal representations are used to compute similarity scores between the target item and each historically interacted item. These scores are discretized and mapped into learnable embeddings as input to the DTA module, while SH (Similarity Histogram) operates directly on the raw similarity scores. Mod-Centric and Mod-Enriched denote the Modality-Centric and Modality-Enriched modeling strategies, respectively. The resulting user interest representations are fused by the CMM module and combined with the target item for CTR prediction.
III-B Target-aware Attention
User interest modeling often employs target-aware attention mechanisms to adaptively capture relevant historical interactions with respect to the target item [5, 13, 15, 12]. In this framework, for a given user interaction sequence of length $L$, let $\mathbf{s}_{j}\in\mathbb{R}^{d}$ denote the embedding of the $j$-th historically interacted item for $j=1,\dots,L$. Let $\mathbf{q}\in\mathbb{R}^{d_{q}}$ denote the embedding of the target item. The sequence embeddings are used to construct the key matrix $\mathbf{K}\in\mathbb{R}^{L\times d_{k}}$ and value matrix $\mathbf{V}\in\mathbb{R}^{L\times d_{v}}$, while the target embedding serves as the query $\mathbf{Q}\in\mathbb{R}^{1\times d_{q}}$. In mini-batch training, these are extended to $\mathbf{K}\in\mathbb{R}^{B\times L\times d_{k}}$, $\mathbf{V}\in\mathbb{R}^{B\times L\times d_{v}}$, and $\mathbf{Q}\in\mathbb{R}^{B\times d_{q}}$, where $B$ is the batch size, $L$ is the maximum sequence length, and $d_{k}$, $d_{v}$, $d_{q}$ are the dimensions of key, value, and query vectors, respectively.
For the $i$-th sample in the batch, let $\mathbf{q}_{i}\in\mathbb{R}^{d}$ denote the query vector corresponding to the target item. Let $\mathbf{s}_{j}\in\mathbb{R}^{d}$ denote the embedding of the $j$-th historically interacted item for $j=1,\dots,L$. The historical embedding matrix is defined as $\mathbf{S}=[\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{L}]^{\top}\in\mathbb{R}^{L\times d}$. The attention mechanism computes a user interest representation $\mathbf{u}_{i}\in\mathbb{R}^{d_{v}}$ as a weighted sum over the value vectors:
$\mathbf{u}_{i}=\sum_{j=1}^{L}\alpha_{ij}\mathbf{v}_{j},$ (3)
where $\mathbf{v}_{j}=\mathbf{V}_{i,j,:}\in\mathbb{R}^{d_{v}}$ denotes the $j$-th value vector in the $i$-th sample, and $\alpha_{ij}$ is the attention weight assigned to the $j$-th item. The weights are computed via softmax normalization over unnormalized alignment scores $e_{ij}\in\mathbb{R}$:
$\alpha_{ij}=\frac{\exp(e_{ij}/\tau)}{\sum_{k=1}^{L}\exp(e_{ik}/\tau)},$ (4)
where $\tau$ is a temperature parameter [32].
DIN [5] and MHTA [13] are the two most representative target-aware attention approaches. The main difference between them is how the score $e_{ij}$ is obtained. In DIN [5], the key and value vectors are set to the input embeddings of the historically interacted items:
$\mathbf{k}_{j}=\mathbf{s}_{j},\quad\mathbf{v}_{j}=\mathbf{s}_{j},$ (5)
and $e_{ij}$ is computed as:
$e_{ij}=\mathrm{MLP}\left([\mathbf{q}_{i};\mathbf{s}_{j};\mathbf{q}_{i}\odot\mathbf{s}_{j};\mathbf{q}_{i}-\mathbf{s}_{j}]\right),$ (6)
where $\cdot$$\cdot$] denotes vector concatenation, $\odot$ denotes element-wise multiplication, and $\mathrm{MLP}$ is a multilayer perceptron with PReLU [33] activations. This design explicitly models both similarity and dissimilarity between the target item and historical interactions.
MHTA [13] computes keys and values via linear projections. Let $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d_{h}\times d}$ be learnable weight matrices, where $d_{h}$ is the attention hidden dimension, which corresponds to the key and value dimensions, i.e., $d_{k}=d_{v}=d_{h}$. The projected key and value vectors are:
$\mathbf{k}_{j}$ $=\mathbf{W}_{k}\mathbf{s}_{j}\in\mathbb{R}^{d_{h}},$ (7)
$\mathbf{v}_{j}$ $=\mathbf{W}_{v}\mathbf{s}_{j}\in\mathbb{R}^{d_{h}},$ (8)
and $e_{ij}$ is computed using scaled dot-product attention [32]:
$e_{ij}=\frac{(\mathbf{W}_{q}\mathbf{q}_{i})^{\top}\mathbf{k}_{j}}{\sqrt{d_{h}}},$ (9)
where $\mathbf{W}_{q}\mathbf{q}_{i}\in\mathbb{R}^{d_{h}}$ serves as the query vector.
Although both approaches follow the same attention paradigm, they differ in the construction of $\mathbf{K}$, $\mathbf{V}$, and the computation of $e_{ij}$. In DIN, $e_{ij}$ depends nonlinearly on both $\mathbf{q}_{i}$ and $\mathbf{s}_{j}$, making it expressive but preventing reuse of intermediate representations across different target items. As a result, the entire attention module must be recomputed for each target item during inference.
In MHTA, the key and value matrices $\mathbf{K}=\mathbf{S}\mathbf{W}_{k}^{\top}$ and $\mathbf{V}=\mathbf{S}\mathbf{W}_{v}^{\top}$ are functions of the historical sequence only and do not depend on the target item. As a result, once computed for a given user sequence, they can be reused across multiple target items without recomputation. This structural property reduces redundant computation and improves inference efficiency in industrial recommendation systems.
The computational bottleneck in DIN arises from the MLP applied to each query-key pair, while in MHTA it stems from the linear projections of the key and value vectors [15, 13]. Our proposed Decoupled Target Attention (DTA), introduced in Section IV-C, builds upon this foundation to support expressive multimodal fusion while preserving inference efficiency.
IV Methodology